# Localized Schrödinger Bridge Sampler

### v1

@@ 1 Step 1: Construct transition Sinkhorn weights $v_{[\alpha]}$
3       construct localised data $\mathcal{X}_{[\alpha]}$;
4       construct kernel matrix $T_{\alpha}\in\mathbf{R}^{M\times M}$ from localised data;
@@ 14                   $X_{[\alpha]}(n)\leftarrow X(n)$ ;
15                   $\Xi_{[\alpha]}(n)\sim{\rm N}(0,\epsilon I_{d_{\alpha}})$;
                   $X_{[\alpha]}(n+1/2)=X_{[\alpha]}(n)+\sqrt{2}\,\Xi_{[\alpha]}(n)$ ;
@@ 23      $x^{(j)}_{\rm s}\leftarrow X(n_{c})$;
Algorithm 1 Localised Schrödinger bridge sampler
@@
We have assumed here a strong form of conditional independence by requesting that (32) holds for all $\epsilon>0$. In general, such a condition will be satisfied approximately for sufficiently small $\epsilon$ only. Compare the EM sampler (16), which provides an accurate approximation to the true transition densities $p(x^{\prime}|x;\epsilon)$ of the underlying diffusion process for $\epsilon>0$ sufficiently small by ignoring higher-order dependencies. In practice, this requires a careful choice of the dependency sets $\Lambda(\alpha)$ which define $x_{[\alpha]}\in\mathbb{R}^{d_{\alpha}}$.
@@ 3.3. Algorithmic properties
We briefly discuss a few important results on the stability and the ergodicity of the proposed localized Langevin samplers, which they essentially inherit from the unlocalized Schrödinger bridge sampler [5].
The following lemma establishes that, since each $w_{\alpha}(x_{[\alpha]})$, $\alpha=1,\ldots,d$, is a probability vector for any $\epsilon>0$, the localized update step (29) is stable. In order to simplify notations, we denote by $m_{\rm loc}(x;\epsilon)\in\mathbb{R}^{d}$ the vector of localized expectation values with components $m_{\alpha}(x_{[\alpha]})$, $\alpha=1,\ldots,d$, defined by (27).
@@
Lemma 1 also establishes stability of the general Langevin sampler defined by (5) with localized $m_{\rm loc}(x;\epsilon)$, i.e.
(39) $X(n+1)=m_{\rm loc}(X(n);\epsilon)+\sqrt{2}\,\Xi(n),\qquad\Xi(n)\sim{\rm N}(0,\epsilon I),$
for all step sizes $\epsilon>0$. Note that $X(n+1)$ is no longer in the convex hull of the data as the original unlocalized Schrödinger bridge (cf. (14)b), but instead is confined to $\mathcal{C}_{M}$ in expectation. The next lemma shows that this Langevin sampler is also geometrically ergodic.
@@
Consider the Lyapunov function $V(x)=1+\|x\|^{2}$ and introduce the set
(40) $\mathcal{C}=\{x:\|x\|\leq R\}$
for suitable $R>0$. Since $m_{\rm loc}(x;\epsilon)\in\mathcal{C}_{M}$ and $\pi$ has compact support, one can find a radius $R>0$, which is independent of the training data $\mathcal{X}$, such that $\mathcal{C}_{M}\subset\mathcal{C}$ and
(41) $\mathbb{E}[V(X(n+1)|X(n)]\leq\lambda V(X(n))$
for all $X(n)\notin\mathcal{C}$ with $10\leq\lambda<$ 1. Furthermore, because of the additive Gaussian noise in (39), there is a probability density function $\rho(x)>0$ on $\mathcal{C}$ and a constant $\delta>0$ such that
(42) ${\rm n}(x^{\prime};m_{\rm loc}(x;\epsilon),2\epsilon I)\geq\delta\rho(x^{\prime})$
for all $x,x^{\prime}\in\mathcal{C}$. Here ${\rm n}(x;m,\Sigma)$ denotes the Gaussian probability density function with mean $m$ and covariance matrix $\Sigma$. In other words, $\mathcal{C}$ is a small set in the sense of [12]. Geometric ergodicity follows from Theorem 15.0.1 in [12]. See also the self-contained presentation in [11]. ∎
@@
We emphasize that, contrary to the unlocalized Schrödinger bridge sampler, the localized $m_{\rm loc}(x;\epsilon)$ is not restricted to the linear subspace spanned by the training data $\mathcal{X}$ in case $M<d$. The localized sampler shares this desirable property for very high-dimensional problems with the localized EnKF [16, 2, 3].
@@
4. Conditional localised Schrödinger bridge sampler
@@
As for the standard Schrödinger bridge sampler [5] the localized sampler lends itself to conditional sampling. Consider samples $x^{(j)}=(z^{(j)},\psi^{(j)})^{T}$ for $j=1,\dots,M$ with $\psi=\psi(z)$. The localized Schrödinger bridge sampler described in Section 3 and Algorithm 1 allows us to learn the joint probability $p(z,\psi)$. To draw samples from the conditional probability $p(\psi|z)$ we may use the localized conditional probability vector $t_{\alpha}(x_{[\alpha]})$ and the Sinkhorn weights $v_{\alpha}$ obtained from the samples $x^{(j)}$ , i.e.  executing lines 1-6 in Algorithm 1. Conditional sampling is achieved by ensuring that at each sampling step the $z$-component of the generated samples $X$ are set to the value $z^{\ast}$ on which we wish to condition. To achieve this we add the conditioning assignment, $Z(n+1)\leftarrow z^{\ast}$, between lines 20 and 21 of Algorithm 1. We demonstrate the performance of the conditional sampler for the multiscale Lorenz-96 model in the next subsection.
@@ 4.1. Conditional sampling for a closure problem
We apply the conditional localized Schrödinger bridge sampler to the multiscale Lorenz-96 model for $K$ slow variables $z_{k}$ which are each coupled to $J$ fast variables $y_{j,k}$ with
@@
(43b) $\frac{d}{dt}y_{j,k}$ $=-cby_{j+1,k}(y_{j+2,k}-y_{j-1,k})-cy_{j,k}+\frac{hc}{b}z_{k}.$
with periodic boundary conditions with $z_{k+K}=z_{k}$, $y_{j,k+K}=y_{j,k}$ and $y_{j+J,k}=y_{j,k+1}$. This $d=K(J+1)$-dimensional model was introduced as a caricature for the mid-latitude atmospheric dynamics [9]. The degree of time-scale separation is controlled by the parameter $c$. The ratio of the amplitudes of the large-scale variables $z_{k}$ and the small-scale variables $y_{j,k}$ is controlled by the parameter $b$. The slow and fast dynamics are coupled with coupling strength $h$. The parameter $F$ denotes external forcing. As equation parameters we choose $K=12$ and $J=24$, i.e. $d=300$, and $F=20$, $c=b=10$ and $h=1$ as in [20, 1, 4]. These parameters lead to chaotic dynamics with a maximal Lyapunov exponent of $\lambda_{\rm{max}}\approx 18.29$ in which the fast variables experience temporal fluctuations which are $10$ times faster and $10$ times smaller than those of the slow variables. This corresponds to the regime of strong coupling in which the dynamics is driven by the fast sub-system [6].
In the climate science and other disciplines one is typically only interested in the slow large-scale dynamics. A direct simulation of the multiscale system (43), however, requires a small time step adapted to the fastest occurring time scale, making long term integration to resolve the slow dynamics computationally infeasible. Scientists hence aim to design a computationally tractable model for the slow variables only in which the effect of the fast dynamics is parameterized. This is the so called closure or subgrid-scale parameterization problem. In particular, we seek a model of the form
@@
with $z=(z_{1},z_{2},\ldots,z_{K})$, $G_{k}(z)=-z_{k-1}(z_{k-2}-z_{k+1})-z_{k}+F$. We assume here that scientists have prior physics-based knowledge about the resolved vector field $G_{k}(z)$ but lack knowledge of the closure term $\psi_{k}(z)$ which parametrizes the effect of the fast unresolved dynamics. The closure term may be deterministic or stochastic, depending on the choice of equation parameters in (43). We will employ the localized Schrödinger bridge sampler to generate samples of the closure term $\psi(z)$ conditioned on the current model state $z(t)$. The sampler will be trained on $M$ samples $x^{(j)}=(z^{(j)},\psi^{(j)})$, $j=1,\ldots,M$.
We obtain $M=40,000$ samples $z^{(j)}\in\mathbb{R}^{K}$ by integrating (43) using a fourth-order Runge-Kutta method with a fixed time step $\Delta t=5\times{10^{-4}}$. Samples of the closure term $\psi^{(j)}\in\mathbb{R}^{K}$ are then determined from the samples $z^{(j)}$ via
@@
for $j=1,\cdots,M$. This defines our $M$ samples $x^{(j)}=(z^{(j)},\psi^{(j)})\in\mathbb{R}^{2K}$ for $j=1,\ldots,M$ to be used to train the Schrödinger bridge sampler.
@@
To numerically integrate the closure model (44) using an Euler discretization
(46) $z(m+1)=z(m)+(G(z(m))+\psi(m))\,\delta t$
@@
in terms of the state vector $z\in\mathbb{R}^{K}$ with a time step $\delta t=10\Delta t$, we generate at each time step $m\geq 0$ a sample $\psi(m)$ conditioned on the current state $z(m)$. These samples should be uncorrelated to the samples drawn at the previous time step. This is achieved by running the localized Schrödinger bridge sampler conditioned on $z^{\ast}=z(m)$ at each time step $m$ for $n_{c}=100$ decorrelation steps (cf. Algorithm 1).
@@
For the localized Schrödinger bridge sampler we employ a time step of $\epsilon=0.1$ and consider a nearest neighbour localisation with $\Lambda(\alpha)=\{\alpha-1,\alpha,\alpha+1,\alpha,K+\alpha-1,K+\alpha,K+\alpha+1\}$ with the obvious periodic extensions for $\alpha=1$ and $\alpha=d$. To account for the varying ranges of $z$ and $\psi$ when estimating the matrices (23) and (25) for fixed time step $\epsilon$, we replace the standard Euclidean product with a scaled one where we divide in the inner product the $z$-variables by $\sigma_{z}$ and the $\psi$-variables by $\sigma_{\psi}$, where $\sigma_{\psi}^{2}$ and $\sigma_{x}^{2}$ denote the climatic variances of the slow variables and the closure term, respectively, estimated from the samples $x^{(j)}$.
@@
Figure 2 shows a comparison of the outputs of the localized Schrödinger bridge sampler with data obtained from simulating the full multi-scale Lorenz-96 system (43). We show results for the covariance of the slow variables $z$, obtained from the samples $z^{(j)}$ of the full multi-scale Lorenz 96 system (43), and of the discretization of the closure scheme (46) over $40,000$ time steps with time step $\delta t$. Samples $\psi(m)$ are generated at each time step using the conditional sampling algorithm. In particular, we show the entries of the rows of the empirical covariance matrix, centred about $k=6$ employing periodicity of the system. It is seen that our localized sampler reproduces the covariance structure of the full system very well. We further show a comparison of the empirical histograms obtained by integrating the closure model (46) with the original samples $z^{(j)}$ which were obtained from a simulation of the full multi-scale Lorenz 96 system (43). The learned non-stiff stochastic closure model (46) is able to reproduce the actual histogram well.
@@
Figure 2. Comparison of the samples obtained from the localized Schrödinger bridge sampler and given samples drawn from the multi-scale Lorenz-96 system (43) using nearest neighbour localization with $\Lambda(\alpha)=\{\alpha-1,\alpha,\alpha+1\,K+\alpha-1,K+\alpha,K+\alpha+1\}$ with the obvious periodic extensions for $\alpha=1$ and $\alpha=d$. We consider $40,000$ new and given samples. Left: Centred rows of the empirical covariance matrix for $z$. The magenta line denotes the mean over all rows. The blue markers denote the empirical covariance for the given samples. Right: Empirical histograms. The closure model used at ten times larger time step.

### v3

@@ 1 Step 1: Construct transition Sinkhorn weights $v_{[\alpha]}$
3       construct localized data $\mathcal{X}_{[\alpha]}$;
4       construct kernel matrix $T_{\alpha}\in\mathbf{R}^{M\times M}$ from localized data;
@@ 15                   $X_{[\alpha]}(n)\leftarrow X(n)$ ;
16                   $\Xi_{[\alpha]}(n)\leftarrow\Xi(n)$ ;
                   $X_{[\alpha]}(n+1/2)=X_{[\alpha]}(n)+\sqrt{2\epsilon}\,\Xi_{[\alpha]}(n)$ ;
@@ 24      $x^{(j)}_{\rm s}\leftarrow X(n_{c})$;
Algorithm 1 Localized Schrödinger bridge sampler
@@
We have assumed here a strong form of conditional independence by requesting that (61) holds for all $\epsilon>0$. In general, such a condition will be satisfied approximately for sufficiently small $\epsilon$ only. Compare the EM sampler (40), which provides an accurate approximation to the true transition densities $p(x^{\prime}|x;\epsilon)$ of the underlying diffusion process for $\epsilon>0$ sufficiently small by ignoring higher-order dependencies. In practice, this requires a careful choice of the dependency set $\Lambda(\alpha)$ which defines $x_{[\alpha]}\in\mathbb{R}^{d_{\alpha}}$.
@@ 3.3. Algorithmic properties
We briefly discuss a few important results on the stability and ergodicity of the proposed localized Langevin samplers, which they essentially inherit from the unlocalized Schrödinger bridge sampler [10].
The following lemma establishes that, since each $w_{\alpha}(x_{[\alpha]})$, $\alpha=1,\ldots,d$, is a probability vector for any $\epsilon>0$, the localized update step (56b) is stable. In order to simplify notations, we denote by $m_{\rm loc}(x;\epsilon)\in\mathbb{R}^{d}$ the vector of localized expectation values with components $m_{\alpha}(x_{[\alpha]})$, $\alpha=1,\ldots,d$, defined by (52).
@@
Lemma 1 also establishes stability of the general Langevin sampler defined by (5) with localized $m_{\rm loc}(x;\epsilon)$ and $S(x;\epsilon)=2\epsilon I$, i.e.,
(68) $X(n+1)=m_{\rm loc}(X(n);\epsilon)+\sqrt{2\epsilon}\,\Xi(n),\qquad\Xi(n)\sim{\rm N}(0,I),$
for all step sizes $\epsilon>0$. Note that $X(n+1)$ is no longer in the convex hull of the data as the original unlocalized Schrödinger bridge (cf. (19)b), but instead is confined to $\mathcal{C}_{M}$ in expectation. The exact gradient structure of the conditional expectation value (24) does no longer hold for the localized $m_{\rm loc}(x;\epsilon)$. The next lemma shows that the localized sampler (68) remains geometrically ergodic.
@@
Consider the Lyapunov function $V(x)=\|x\|^{2}$ and introduce the ball
(69) $\mathcal{B}_{R}=\{x\in\mathbb{R}^{d}:\|x\|\leq R\}$
of radius $R>0$. Since $m_{\rm loc}(x;\epsilon)\in\mathcal{C}_{M}$ and $\pi$ has compact support, one can find a radius $R>0$, which is independent of the training data $\mathcal{X}$, such that $\mathcal{C}_{M}\subset\mathcal{B}_{R}$ and
(70) $\mathbb{E}[V(X(n+1))|X(n)]\leq\lambda V(X(n))$
for all $X(n)\notin\mathcal{B}_{R}$ with $10\leq\lambda<$ 1. Furthermore, because of the additive Gaussian noise in (68), there is a constant $\delta>0$ such that
(71) ${\rm n}(x^{\prime};m_{\rm loc}(x;\epsilon),2\epsilon I)\geq\delta$
for all $x,x^{\prime}\in\mathcal{B}_{R}$. Here ${\rm n}(x;m,C)$ denotes the Gaussian probability density function with mean $m$ and covariance matrix $C$. In other words, $\mathcal{B}_{R}$ is a small set in the sense of [19]. Geometric ergodicity follows from Theorem 15.0.1 in [19]. See also the self-contained presentation in [18]. ∎
@@
We emphasize that, contrary to the unlocalized Schrödinger bridge sampler, the localized $m_{\rm loc}(x;\epsilon)$ is not restricted to the linear subspace of $\mathbb{R}^{d}$ spanned by the training data $\mathcal{X}\in\mathbb{R}^{d\times M}$ in case $M<d$. The localized sampler shares this desirable property with the localized EnKF [25, 4, 7].
3.4. Localized kernel-denoising
Localizing the KDE-based denoiser (35) follows along the same lines. We first note that a component-wise formulation of Tweedie’s formula (31) leads to
(72) $\partial_{x_{\alpha}}\log\pi_{\epsilon}(x)=-\frac{1}{\epsilon}\left(x_{\alpha}-\mathbb{E}[x_{\alpha}^{\prime}|x]\right),$
$\alpha=1,\ldots,d$. Upon assuming the conditional independence relation
(73) $\mathbb{E}[x_{\alpha}^{\prime}|x]=\mathbb{E}[x_{\alpha}^{\prime}|x_{[\alpha]}]$
one finds that
(74) $\partial_{x_{\alpha}}\log\pi_{\epsilon}(x)=\partial_{x_{\alpha}}\log\pi_{\epsilon}(x_{[\alpha]}).$
We approximate the restricted density $\pi_{\epsilon}(x_{[\alpha]})$ by the localized KDE estimator
(75) $\tilde{\pi}_{\epsilon}(x_{[\alpha]})\propto\sum_{j=1}^{M}\exp\left(-\frac{1}{2\epsilon}\|x_{[\alpha]}-x_{[\alpha]}^{(j)}\|^{2}\right),$
which results in
(76) $\partial_{x_{\alpha}}\log\pi_{\epsilon}(x)\approx-\frac{1}{\epsilon}\left(x_{\alpha}-\sum_{j=1}^{M}x_{\alpha}^{(j)}\,{\tilde{w}}_{\alpha}^{(j)}(x_{[\alpha]})\right)$
with localized weights
(77) $\tilde{w}_{\alpha}^{(j)}(x_{[\alpha]}):=\frac{\exp\left(-\frac{1}{2\epsilon}\|x_{[\alpha]}-x_{[\alpha]}^{(j)}\|^{2}\right)}{\sum_{j=1}^{M}\exp\left(-\frac{1}{2\epsilon}\|x_{[\alpha]}-x_{[\alpha]}^{(j)}\|^{2}\right)}$
for $j=1,\ldots,M$. We collect these weights in the vector $\tilde{w}_{\alpha}(x;\epsilon)\in\mathbb{R}^{M}$. One finally obtains the following localized KDE update step for $X_{\alpha}(n+1/2)$:
(78) $X_{\alpha}(n+1)=D_{\alpha}(X(n+1/2);\epsilon):=\mathcal{X}_{\alpha}\,\tilde{w}_{\alpha}(X_{[\alpha]}(n+1/2)),\qquad\alpha=1,\ldots,d,$
which can be employed whenever (73) holds to sufficient accuracy.
@@ 3.4. Localized kernel-denoising
4. Localised Schrödinger bridge sampler for temporal stochastic processes
In this section, we consider temporal stochastic processes $Z(t_{k})\in\mathbb{R}^{s}$ with $t_{k}=k\Delta t$ and $k=0,\ldots,K$, and assume that $M$ realizations
(79) $x^{(j)}=\{Z^{(j)}(t_{k})\}_{k=0}^{K}\in\mathbb{R}^{d},\qquad d=(K+1)s,$
$j=1,\ldots,M$, of such a process have become available. We furthermore assume that the generating process is Markovian, i.e., $Z(t_{k+1})$ is conditionally independent of all $Z(t_{l})$ with $l<k$ and $l>k+1$. Such a setting provides a perfect application of the localization strategy proposed in Section 3. More specifically, we obtain the following subsets for localization in terms of the entries $x_{\alpha}$ of the augmented state vector $x\in\mathbb{R}^{d}$: $\Lambda(\alpha)=\{1,\ldots,s\}$ for $\alpha\in\{1,\ldots,s\}$ and
(80) $\Lambda(\alpha)=\{sl+1,\ldots,s(l+2)\}$
for $\alpha\in\{s(l+1)+1,\ldots,s(l+2)\}$ and $l=0,\ldots,K-2$.
As a numerical illustration we consider the bimodal stochastic differential equation (SDE)
(81) $\frac{d}{dt}Z(t)=-Z(t)^{3}+Z(t)+\sqrt{0.2}\frac{d}{dt}B(t),\qquad Z(0)\sim{\rm N}(0,1),$
where $B(t)$ denotes standard Brownian motion. Here $s=1$ and we sample solutions in time-intervals of length $\Delta t=5$ over $K=100$ intervals; hence the dimension of the augmented state vector $x\in\mathbb{R}^{d}$ becomes $d=101$. The training data consists of $M=1,000$ independent realizations, which were obtained with a small step-size EM algorithm applied to (81).
Results from the localized Schrödinger bridge split-step sampler (56) with constant diffusion $S(x;\epsilon)=2\epsilon I$ with $\epsilon=0.0025$ and $N=25,000$ generated samples can be found in Figures 3 and 4, respectively. The numerical results demonstrate that the localized Schrödinger bridge sampler can successfully generated samples for this rather high-dimensional ($d=101$) and nonlinear problem given only $M=1,000$ training samples.
We also implemented the localized KDE-based denoiser (78). The results are virtually indistinguishable from the results obtained from the localized Schrödinger bridge sampler.
Figure 3. Generated trajectories for the bimodal SDE (81) using the localized Schrödinger bridge split-step sampler with constant diffusion (56). Left panel: $100$ trajectories out of $M=1,000$ training samples; right panel: $100$ trajectories out of $N=25,000$ generated samples. The computed transition rates (relative number of sign changes along trajectories) agree well with $9\%$ for the training data and $11\%$ for the generated data.
Figure 4. Normalized empirical histograms of training and generated data for the bimodal SDE (81) using the localized Schrödinger bridge split-step sampler with constant diffusion (56). We show results over all $1,000$ training and $25,000$ generated data points. The invariant distribution of the bimodal SDE is well reproduced by the generated data; the dispersion of the generated data in each of its two modes being slightly smaller than the one from the training data, which has also been observed for the split-step scheme in Figure 2.
5. Conditional localized Schrödinger bridge sampler
@@
As for the standard Schrödinger bridge sampler [10] the localized sampler lends itself to conditional sampling. Consider samples $x^{(j)}=(z^{(j)},\psi^{(j)})$ for $j=1,\dots,M$. The localized Schrödinger bridge sampler described in Section 3 and Algorithm 1 allows us to learn the joint probability measure $\nu({\rm d}z,{\rm d}\psi)$. To draw samples from the conditional probability measure $\nu({\rm d}\psi|z)$ we may use the localized conditional probability vector $t_{\alpha}(x_{[\alpha]})$ and the Sinkhorn weights $v_{\alpha}$ obtained from the samples $x^{(j)}$ , i.e., executing lines 1-6 in Algorithm 1. Conditional sampling is achieved by ensuring that at each sampling step the $z$-component of the generated samples $X(n)$ are set to the value $z^{\ast}$ on which we wish to condition. To achieve this we add the conditioning assignment, $Z(n)\leftarrow z^{\ast}$, between lines 11 and 12 of Algorithm 1. We demonstrate the performance of the conditional sampler for the multi-scale Lorenz-96 model in the next subsection.
@@ 5.1. Conditional sampling for a closure problem
We apply the conditional localized Schrödinger bridge sampler to the multi-scale Lorenz-96 model for $K$ slow variables $z_{k}$ which are each coupled to $J$ fast variables $y_{j,k}$ and evolve according to
@@
(82b) $\frac{d}{dt}y_{j,k}$ $=-cby_{j+1,k}(y_{j+2,k}-y_{j-1,k})-cy_{j,k}+\frac{hc}{b}z_{k}$
with periodic boundary conditions $z_{k+K}=z_{k}$, $y_{j,k+K}=y_{j,k}$ and $y_{j+J,k}=y_{j,k+1}$. This $d=K(J+1)$-dimensional model was introduced as a caricature for the mid-latitude atmospheric dynamics [16]. The degree of time-scale separation is controlled by the parameter $c$. The ratio of the amplitudes of the large-scale variables $z_{k}$ and the small-scale variables $y_{j,k}$ is controlled by the parameter $b$. The slow and fast dynamics are coupled with coupling strength $h$. The parameter $F$ denotes external forcing. As equation parameters we choose $K=12$ and $J=24$, i.e. $d=300$, and $F=20$, $c=b=10$ and $h=1$ as in [34, 2, 9]. These parameters lead to chaotic dynamics with a maximal Lyapunov exponent of $\lambda_{\rm{max}}\approx 18.29$ in which the fast variables experience temporal fluctuations which are $10$ times faster and $10$ times smaller than those of the slow variables. This corresponds to the regime of strong coupling in which the dynamics is driven by the fast sub-system [11].
In the climate science and other disciplines one is typically only interested in the slow large-scale dynamics. A direct simulation of the multi-scale system (82), however, requires a small time step adapted to the fastest occurring time scale, making long term integration to resolve the slow dynamics computationally infeasible. Scientists hence aim to design a computationally tractable model for the slow variables only in which the effect of the fast dynamics is parameterized. This is the so called closure or subgrid-scale parameterization problem. In particular, we seek a model of the form
@@
for $z=(z_{1},z_{2},\ldots,z_{K})$ with $G_{k}(z)=-z_{k-1}(z_{k-2}-z_{k+1})-z_{k}+F$. We assume that scientists have prior physics-based knowledge about the resolved vector field $G_{k}(z)$ but lack knowledge of the closure term $\psi_{k}(z)$ which parametrizes the effect of the fast unresolved dynamics. The closure term may be deterministic or stochastic, depending on the choice of equation parameters in (82). We will employ the localized Schrödinger bridge sampler to generate samples of the closure term $\psi(z)$ conditioned on the current model state $z(t)$. The sampler will be trained on $M$ samples $x^{(j)}=(z^{(j)},\psi^{(j)})$, $j=1,\ldots,M$, which consists of a time series with $x^{(j)}=x(t_{j})$ with $t_{j}=j\Delta t$ and $\Delta t=5\times{10^{-3}}$.
We obtain $M=40,000$ samples $z^{(j)}\in\mathbb{R}^{K}$ by integrating (82) using a fourth-order Runge–Kutta method with a fixed time step $\delta t=5\times{10^{-4}}$ and collecting the state in time intervals of $\Delta t=10\,\delta t$. Samples of the closure term $\psi^{(j)}\in\mathbb{R}^{K}$ are then determined from the samples $z^{(j)}$ via
@@
for $j=1,\cdots,M-1$. This defines $M-1$ samples $x^{(j)}=(z^{(j)},\psi^{(j)})\in\mathbb{R}^{2K}$ for $j=1,\ldots,M-1$ to be used to train the Schrödinger bridge sampler.
@@
To numerically integrate the closure model (83) in terms of the state vector $z\in\mathbb{R}^{K}$ we employ an Euler discretization
(85) $z(m+1)=z(m)+(G(z(m))+\psi(m|z(m)))\,\Delta t$
@@
with a time step $\Delta t$. At each time step $m\geq 0$ we generate a sample $\psi(m|z(m))$ conditioned on the current state $z(m)$. These samples should be uncorrelated to the samples drawn at the previous time step. This is achieved by running the localized Schrödinger bridge sampler conditioned on $z^{\ast}=z(m)$ at each time step $m$ for $n_{c}=100$ decorrelation steps (cf. Algorithm 1).
@@
For the localized Schrödinger bridge sampler we employ a parameter of $\epsilon=0.1$ and consider a nearest neighbor localization with $\Lambda(\alpha)=\{\alpha-1,\alpha,\alpha+1,\alpha,K+\alpha-1,K+\alpha,K+\alpha+1\}$ with the obvious periodic extensions for $\alpha=1$ and $\alpha=d$. To account for the varying ranges of $z$ and $\psi$ when estimating the matrices (47) and (50) for fixed parameter $\epsilon$, we replace the standard Euclidean product with a scaled one where we divide the inner product in the $z$-variables by $\sigma_{z}$ and the $\psi$-variables by $\sigma_{\psi}$, where $\sigma_{\psi}^{2}$ and $\sigma_{x}^{2}$ denote the climatic variances of the slow variables and the closure term, respectively, estimated from the samples $x^{(j)}$.
@@
Figures 5 and 6 show a comparison of the outputs of the localized Schrödinger bridge sampler with data obtained from simulating the full multi-scale Lorenz-96 system (82). We show results for the covariance of the slow variables $z$, obtained from the samples $z^{(j)}$ of the full multi-scale Lorenz-96 system (82), and of the discretization of the closure scheme (85). We show in Figure 5 a comparison of the empirical histograms of $z$ obtained by integrating the closure model (85) with the original samples $\{z^{(j)}\}_{j=1}^{M}$ which were obtained from a simulation of the full multi-scale Lorenz-96 system (82). The non-stiff trained stochastic closure model (85) is able to reproduce the actual histogram well. We further show a scatter plot of the stochastic closure term $\psi$ obtained from the full Lorenz-96 system (82) and obtained from the localized conditional Schrödinger bridge. The closure term is well represented by the localized conditional Schrödinger bridge. In Figure 6 we show the entries of the rows of the empirical covariance matrix, centered about $k=6$ employing periodicity of the system. It is seen that our localized sampler reproduces the covariance structure of the full system very well. We further show that the temporal autocorrelation structure of the Lorenz-96 system is well reproduced by the localized conditional Schrödinger bridge.
@@
Figure 5. Comparison of the samples obtained from the localized Schrödinger bridge sampler and given samples drawn from the multi-scale Lorenz-96 system (82) using nearest neighbor localization with $\Lambda(\alpha)=\{\alpha-1,\alpha,\alpha+1\,K+\alpha-1,K+\alpha,K+\alpha+1\}$ with the obvious periodic extensions for $\alpha=1$ and $\alpha=d$. We consider $40,000$ new and given samples. Left: Empirical histograms. Right: Scatter plot of the closure term $\psi$ as a function of $z$.
Figure 6. Comparison of the samples obtained from the localized Schrödinger bridge sampler and given samples drawn from the multi-scale Lorenz-96 system (82) using nearest neighbor localization with $\Lambda(\alpha)=\{\alpha-1,\alpha,\alpha+1\,K+\alpha-1,K+\alpha,K+\alpha+1\}$ with the obvious periodic extensions for $\alpha=1$ and $\alpha=d$. We consider $40,000$ new and given samples. Left: Centered rows of the empirical covariance matrix for $z$. The magenta line denotes the mean over all rows. The blue markers denote the empirical covariance for the given samples. Right: Autocovariance function $R(\tau)$.